# SemiOccam: A Robust Semi-Supervised Image Recognition Network Using Sparse Labels

### v2

@@ Abstract
We present SemiOccam, an image recognition network that leverages semi-supervised learning in a highly efficient manner. Existing works often rely on complex training techniques and architectures, requiring hundreds of GPU hours for training, while their generalization ability when dealing with extremely limited labeled data remains to be improved. To address these limitations, we construct a hierarchical mixture density classification decision mechanism by optimizing mutual information between feature representations and target classes, compressing redundant information while retaining crucial discriminative components. Experimental results demonstrate that our method achieves state-of-the-art performance on various datasets when using negligible labeled samples, and its simple architecture keeps training time to minute-level. Notably, this paper reveals a long-overlooked data leakage issue in the STL-10 dataset for semi-supervised learning tasks and removes duplicates to ensure the reliability of experimental results. We also release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research in future semi-supervised learning.  [^1] Code available at https://github.com/Shu1L0n9/SemiOccam.  [^2] CleanSTL-10 available at https://huggingface.co/datasets/Shu1L0n9/CleanSTL-10.
@@ 1 Introduction
Deep learning has achieved remarkable success in image classification tasks. However, its performance is highly dependent on large-scale annotated datasets. In real-world applications, labeled data is often scarce and expensive to acquire. Therefore, semi-supervised learning (SSL) has received widespread attention in recent years as an effective solution. The goal of semi-supervised learning is to train models using a small amount of labeled data and large amount of unlabeled data, thereby improving the model’s generalization ability. Nevertheless, most existing SSL methods, such as consistency regularization and pseudo-labeling, rely on complex training strategies and network designs, leading to training costs of hundreds of GPU hours even on high-performance computing resources like NVIDIA V100, and their generalization ability when labeled data are extremely limited still needs improvement. Moreover, although some recent works have explored Vision Transformers (ViT) in SSL, they have yet to fully exploit its advantages.
Gaussian Mixture Models (GMMs) are powerful probabilistic models that assume data is generated from a mixture of multiple Gaussian distributions. The inspiration for this research comes from Miller & Uyar’s systematic study of the “Mixture of Experts (MoE)” structure based on the joint probability model of features and labels, as well as the principle of Occam’s razor.
@@
According to Occam’s razor, among competing hypotheses that explain the data well, the simplest one is preferred. This motivates us to design a model that balances complexity and interpretability, ensuring efficient and effective feature representation without unnecessary overfitting. In our research, we found that the performance of GMMs is highly related to the quality of feature vectors. If the features lack representativeness or have insufficient discriminative power, the classification performance will be severely affected.
@@
To address this, we designed a feature encoding framework based on a self-attention mechanism to fully exploit the global contextual information in images and generate more discriminative feature representations. Simultaneously, we incorporated semi-supervised learning strategies, utilizing a pseudo-labeling mechanism to efficiently leverage unlabeled data.
@@
This approach achieves superior classification performance within the framework of mixture density estimation, significantly overcomes the limitations of traditional methods in terms of feature representation capability, training cost and model generalization performance. Taking the Dino with ViT-large backbone as our feature extraction model, we validated the proposed semi-supervised Gaussian mixture model (SGMM) method on various datasets, results show that our method achieves state-of-the-art performance within minutes on a single Tesla T4 GPU and exhibits good stability and robustness, with strong potential for edge deployment.
Figure 1: Performance Comparison. Comparison of our method with recent semi-supervised learning models on datasets with different numbers of labels, achieving state-of-the-art performance with low training cost.
@@
  • We design a simple but performant semi-supervised learning strategy that effectively leverages negligible labels, achieves state-of-the-art performance.
@@ 2.1 Semi-Supervised Learning and Gaussian Mixture Models
Pseudo-Labeling is one of the classic methods in SSL, with the core idea of using an existing model to predict unlabeled data Yarowsky (1995) and adding its predicted labels to the training set as pseudo-labels. Representative works include Lee et al. (2013) and Zhang et al. (2021), which iteratively enhance the labeled dataset through self-training. Although pseudo-labeling methods can effectively improve model performance, they are often affected by the quality of pseudo-labels and can easily lead to model overfitting. The study by Mishra et al. (2024) explores the issue of overconfidence caused by pseudo-labels in semi-supervised learning, analyzes the source of calibration errors, and aims to improve the reliability of the model by evaluating and potentially improving methods.
Consistency Regularization methods improve the generalization ability of models by forcing them to maintain consistency on unlabeled data. The Virtual Adversarial Training (VAT) method proposed by Miyato et al. (2018) enhances the robustness of the model by perturbing the input data and maintaining the stability of the prediction results. The Mean Teacher method proposed by Tarvainen & Valpola (2017) also further improves model performance by using the consistency between the predictions of the teacher model and the output of the student model. Although these methods have achieved good results in many tasks, they often rely on complex training strategies and architectural designs, and their generalization ability across different datasets still needs to be improved.
Semi-Supervised Gaussian Mixture Model is the basic probabilistic framework of our method. It combines GMM with unlabeled data and proposes a learning method that simultaneously considers both labeled and unlabeled data by maximizing the joint log-likelihood function. Our method is inspired by the classifier structure and learning algorithm proposed by Miller & Uyar, which effectively leverages unlabeled data to improve performance. In addition, many impressive works, such as Zong et al. (2018), have extended GMMs by incorporating deep autoencoders for tasks like anomaly detection.
Most related to ours is the model of Zhao et al. (2023), which also uses ViT for image embedding and then applies GMM for classification. This model is very similar to SemiOccam, but it addresses the Generalized Category Discovery (GCD) problem from Vaze et al. (2022), aiming to discover new categories, whereas our work goes further to demonstrate that combining deep learning with SGMM can outperform state-of-the-art methods, especially in scenarios with extremely limited labeled data, yielding significant advantages.
@@ 2.2 Deep neural networks and Visual Transformers
Deep Neural Networks have made significant progress in various supervised learning tasks in recent years. The breakthrough achieved by AlexNet, proposed by Krizhevsky et al. (2012), in image classification tasks laid the foundation for the development of deep learning. To improve the generalization ability of Deep Neural Networks (DNNs), many methods have been proposed to address overfitting issues, such as Dropout by Srivastava et al. (2014) and BatchNormalization by Ioffe & Szegedy (2015). The Residual Network (ResNet) proposed by He et al. (2016) effectively solved the gradient vanishing problem in deep network training by introducing residual connections, making it possible to train deeper networks. Subsequently, DLA by Yu et al. (2018) introduced a dynamic hierarchical attention mechanism on top of ResNet, further optimizing layer-wise interaction and feature fusion.
Visual Transformers are image classification models based on the Transformer architecture that capture global information in images through self-attention mechanisms. Dosovitskiy et al. (2021) proposed dividing images into fixed-size patches and applying self-attention mechanisms. In this way, ViT demonstrates strong performance in image classification tasks, especially on large-scale datasets, exhibiting higher generalization ability compared to traditional CNN architectures. The success of ViT has inspired many researchers to try applying it to self-supervised learning. The DINO method proposed by Caron et al. (2021) learns useful feature representations through self-supervised pre-training of ViT models. These representations can be fine-tuned on a small amount of labeled data to achieve high classification accuracy.
@@ 2.3 Self-Supervised Learning and Data Leakage Issues
Self-Supervised Learning is an unsupervised learning approach that automatically learns feature representations through pretext tasks (e.g., image jigsaw puzzles, predicting future frames). The CLIP model proposed by Radford et al. (2021) achieves strong image understanding capabilities through self-supervised learning of image-text contrast, making its performance on unlabeled data very prominent.
In addition, methods like SimCLR from Chen et al. (2020) and MoCo from He et al. (2020) train neural networks through contrastive learning, enabling them to learn discriminative features without labeled data. These methods provide higher quality representations for feature extractors, thereby improving the performance of downstream tasks (such as classification, face recognition, etc.).
Data leakage issue has always been a potential challenge in machine learning. Off-topic images, near-duplicate images, and labeling errors in benchmark datasets can lead to inaccurate estimations of model performance. Handling and removing duplicate samples has become an issue that cannot be ignored in semi-supervised learning. Therefore, Gröger et al. (2024) re-examines the task of data cleaning and formalizes it as a ranking problem.
@@ 3.1 Powerful Feature Extractor
where $\mathbf{Q}=\mathbf{Z}^{(i)}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{Z}^{(i)}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{Z}^{(i)}\mathbf{W}_{V}$. Through multiple layers of Transformer encoding and Principal Component Analysis (PCA) dimensionality reduction, we finally obtain high-quality image feature vectors $\mathbf{f}_{i}\in\mathbb{R}^{d}$ where $d<D$.
@@ 3.2 Semi-supervised Mixture of Experts Classifier
Figure 2: Overview of our SemiOccam network. The network consists of a feature extractor and a semi-supervised mixture of experts classifier. An example of a sub-model is shown in the ellipse at the bottom right, where the circle labeled 1 represents a feature vector with a corresponding label, 2 represents a pseudo-labeled feature vector, and 3 represents an unlabeled feature vector.
The classifier aims to maximize the following joint log-likelihood function:
$\mathcal{L}(\Theta)=\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})}_{\text{unlabeled data likelihood}}+\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})P(c_{i}|l)}_{\text{labeled data likelihood}}$ (3)
where $\Theta=\{\pi_{1},\ldots,\pi_{L},\theta_{1},\ldots,\theta_{L}\}$ is the set of model parameters with $\sum_{l=1}^{L}\pi_{l}=1$, $\pi_{l}$ represents the mixing coefficient of the $l$-th Gaussian component, $p(\mathbf{f}_{i}|\theta_{l})$ is the probability density function of the $l$-th Gaussian component with parameters $\theta_{l}=\{\bm{\mu}_{l},\Sigma_{l}\}$, and $P(c_{i}|l)$ represents the conditional probability that a feature belongs to class $c_{i}$ given Gaussian component $l$.
@@
$\bm{\mu}_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\mathbf{f}_{i}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\mathbf{f}_{i}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (6)
$\Sigma_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}M_{il}^{(t)}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}M_{il}^{(t)}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (7)
where $M_{il}^{(t)}=(\mathbf{f}_{i}-\bm{\mu}_{l}^{(t)})(\mathbf{f}_{i}-\bm{\mu}_{l}^{(t)})^{\top}$.

### v3

@@ Abstract
We present SemiOccam, an image recognition network that leverages semi-supervised learning in a highly efficient manner. Existing works often rely on complex training techniques and architectures, requiring hundreds of GPU hours for training, while their generalization ability when dealing with extremely limited labeled data remains to be improved. To address these limitations, we construct a hierarchical mixture density classification decision mechanism by optimizing mutual information between feature representations and target classes, compressing redundant information while retaining crucial discriminative components. Experimental results demonstrate that our method achieves state-of-the-art performance on three commonly used datasets, with accuracy exceeding 95% on two of them, using only 4 labeled samples per class, and its simple architecture keeps training time to minute-level. Notably, this paper reveals a long-overlooked data leakage issue in the STL-10 dataset for semi-supervised learning tasks and removes duplicates to ensure the reliability of experimental results. We also release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research in future semi-supervised learning.  [^1] Code available at https://github.com/Shu1L0n9/SemiOccam. [^2] CleanSTL-10 available at https://huggingface.co/datasets/Shu1L0n9/CleanSTL-10.
@@ 1 Introduction
Deep learning has achieved remarkable success in image classification tasks. However, its performance is highly dependent on large-scale annotated datasets. In real-world applications, labeled data is often scarce and expensive to acquire. Therefore, semi-supervised learning (SSL) has received widespread attention in recent years as an effective solution. The goal of SSL is to train models using a small amount of labeled data and large amount of unlabeled data, thereby improving the model’s generalization ability. Nevertheless, most existing SSL methods, such as consistency regularization and pseudo-labeling, rely on complex training strategies and network designs, leading to training costs of hundreds of GPU hours [27][39][33] even on high-performance computing resources like NVIDIA V100, and their generalization ability when labeled data are extremely limited still needs improvement. Moreover, although some recent works have explored Vision Transformers (ViT) in SSL [40], they have yet to fully exploit its advantages.
@@
The inspiration for this research stems from the principle of Occam’s razor. This principle posits that among competing hypotheses that explain the data well, the simplest one is preferred. This motivates us to design a solution that balances model complexity and interpretability, ensuring efficient and effective feature representation while preventing unnecessary overfitting.
Aligned with this ethos, we delve into Gaussian Mixture Models (GMM), powerful probabilistic models that assume data is generated from a mixture of multiple Gaussian distributions. This exploration also draws from the systematic study by Miller & Uyar on the ”Mixture of Experts (MoE)” structure based on the joint probability model of features and labels. In our research, we found that the performance of GMM is highly related to the quality of feature vectors. If the features lack representativeness or have insufficient discriminative power, the classification performance will be severely affected.
@@
To address this, we designed a feature encoding framework based on a self-attention mechanism to fully exploit the global contextual information in images and generate more discriminative feature representations. This enhanced feature quality is crucial for optimizing MoE classifier’s performance. Furthermore, to effectively leverage large amounts of unlabeled data, we incorporated a semi-supervised learning strategy utilizing a pseudo-labeling mechanism. An overview of our network, illustrating the integration of these approaches, is provided in Figure 1.
Figure 1: Overview of our SemiOccam network. The network consists of a feature extractor and a semi-supervised mixture of experts classifier. An example of a sub-model is shown in the ellipse at the bottom right, where the circle labeled ① represents a feature vector with a corresponding label, ② represents a pseudo-labeled feature vector, and ③ represents an unlabeled feature vector.
@@
This approach achieves superior classification performance within the framework of mixture density estimation, significantly overcomes the limitations of traditional methods in terms of training cost and model generalization performance. Taking the Dino with ViT-large backbone as our feature extraction model, we validated the proposed semi-supervised gaussian mixture model (SGMM) method on various datasets, results show that our method achieves state-of-the-art performance within minutes on a single Tesla T4 accelerator and exhibits good stability and robustness, enabling fast iteration and potential for edge deployment.
@@
  • We propose a unified framework that effectively exploits unlabeled data by seamlessly integrating deep representation extraction and semi-supervised learning.
@@ 2.1 Semi-Supervised Learning and Gaussian Mixture Models
Pseudo-Labeling is one of the classic methods in SSL, with the core idea of using an existing model to predict unlabeled data [37] and adding its predicted labels to the training set as pseudo-labels. Representative works include [12] and [39], which iteratively enhance the labeled dataset through self-training. Although pseudo-labeling methods can effectively improve model performance, they are often affected by the quality of pseudo-labels and can easily lead to model overfitting. The study by [16] explores the issue of overconfidence caused by pseudo-labels in semi-supervised learning, analyzes the source of calibration errors, and aims to improve the reliability of the model by evaluating and potentially improving methods.
Consistency Regularization methods improve the generalization ability of models by forcing them to maintain consistency on unlabeled data. The Virtual Adversarial Training method proposed by [17] enhances the robustness of the model by perturbing the input data and maintaining the stability of the prediction results. The Mean Teacher method proposed by [30] also further improves model performance by using the consistency between the predictions of the teacher model and the output of the student model. Although these methods have achieved good results in many tasks, they often rely on complex training strategies and architectural designs, and their generalization ability across different datasets still needs to be improved.
Semi-Supervised Gaussian Mixture Model is the basic probabilistic framework of our method. It combines GMM with unlabeled data and proposes a learning method that simultaneously considers both labeled and unlabeled data by maximizing the joint log-likelihood function. Our method is inspired by the classifier structure and learning algorithm proposed by Miller & Uyar, which effectively leverages unlabeled data to improve performance. In addition, many impressive works, such as [42], have extended GMM by incorporating deep autoencoders for tasks like anomaly detection.
Most related to ours is the model of [40], which also uses ViT for image embedding and then applies GMM for classification. This model is very similar to SemiOccam, but it addresses the generalized category discovery problem from [31], aiming to discover new categories, whereas our work goes further to demonstrate that combining deep learning with GMM can outperform state-of-the-art methods, especially in scenarios with extremely limited labeled data, yielding significant advantages.
@@ 2.2 Deep neural networks and Visual Transformers
Deep Neural Networks have made significant progress in various supervised learning tasks in recent years. The breakthrough achieved by AlexNet, proposed by [11], in image classification tasks laid the foundation for the development of deep learning. To improve the generalization ability of Deep Neural Networks (DNN), many methods have been proposed to address overfitting issues, such as Dropout by [28] and Batch Normalization by [9]. The Residual Network (ResNet) proposed by [7] effectively solved the gradient vanishing problem in deep network training by introducing residual connections, making it possible to train deeper networks. Subsequently, DLA by [38] introduced a dynamic hierarchical attention mechanism on top of ResNet, further optimizing layer-wise interaction and feature fusion.
Visual Transformers are image classification models based on the Transformer architecture that capture global information in images through self-attention mechanisms. [5] proposed dividing images into fixed-size patches and applying self-attention mechanisms. In this way, ViT demonstrates strong performance in image classification tasks, especially on large-scale datasets, exhibiting higher generalization ability compared to traditional convolution neural network (CNN) architectures. The success of ViT has inspired many researchers to try applying it to self-supervised learning. The DINO method proposed by [1] learns useful feature representations through self-supervised pre-training of ViT models. These representations can be fine-tuned on a small amount of labeled data to achieve high classification accuracy.
Similar to our work, these two methods integrate ViT with machine learning techniques. [32] combines CNN and ViT to better capture both local and global features, while [25] enhances hyperspectral image classification by refining the attention mechanism and adding an auxiliary branch. Both approaches optimize the ViT architecture to improve performance and efficiency in specific tasks.
@@ 2.3 Self-Supervised Learning and Data Leakage Issues
Self-Supervised Learning is an unsupervised learning approach that automatically learns feature representations through pretext tasks (e.g., image jigsaw puzzles, predicting future frames). The CLIP model proposed by [23] achieves strong image understanding capabilities through self-supervised learning of image-text contrast, making its performance on unlabeled data very prominent.
In addition, methods like SimCLR from [3] and MoCo from [8] train neural networks through contrastive learning, enabling them to learn discriminative features without labeled data. These methods provide higher quality representations for feature extractors, thereby improving the performance of downstream tasks (such as classification, face recognition, etc.).
Data leakage issue has always been a potential challenge in machine learning. Off-topic images, near-duplicate images, and labeling errors in benchmark datasets can lead to inaccurate estimations of model performance. Handling and removing duplicate samples has become an issue that cannot be ignored in semi-supervised learning. Therefore, Gröger et al.’s work re-examines the task of data cleaning and formalizes it as a ranking problem.
@@ 3.1 Powerful Feature Extractor
Figure 2: Feature Extraction Architecture. This diagram illustrates the ViT architecture specifically configured as a feature extractor. Input images are processed through patch embedding and a Transformer Encoder, yielding high-dimensional feature vectors for downstream tasks.
@@
where $\mathbf{Q}=\mathbf{Z}^{(i)}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{Z}^{(i)}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{Z}^{(i)}\mathbf{W}_{V}$. Through multiple layers of Transformer encoding and Principal Component Analysis (PCA) dimensionality reduction, we finally obtain high-quality image feature vectors $\mathbf{f}_{i}\in\mathbb{R}^{d}$ where $d<D$. The overall feature extraction pipeline is illustrated in Figure 2.
@@ 3.2 Semi-supervised Mixture of Experts Classifier
Figure 3: Illustration of our Classifier. The figure visualizes a model with three gaussian components. The solid spheres represent labeled features, while the hollow spheres denote unlabeled features.
To exploit both labeled and unlabeled data, we formulate a semi-supervised maximum likelihood estimation objective. The overall objective is:
$\mathcal{L}(\Theta)=\mathcal{L}_{1}(\Theta)+\lambda\mathcal{L}_{2}(\Theta)$ (3)
where $\mathcal{L}_{1}$ corresponds to the supervised loss, $\mathcal{L}_{2}$ to the unsupervised loss, and $\lambda$ balances the two terms. In our implementation, we set $\lambda=1$ for simplicity.
Specifically,
$\mathcal{L}_{1}(\Theta)=\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\log p_{\Theta}(\mathbf{f}_{i},c_{i}),\qquad\mathcal{L}_{2}(\Theta)=\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\log p_{\Theta}(\mathbf{f}_{i})$ (4)
The joint likelihood of all samples can thus be written as:
$\mathcal{L}(\Theta)=\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})P(c_{i}|l)}_{\text{labeled data likelihood}}+\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})}_{\text{unlabeled data likelihood}}$ (5)
where $\Theta=\{\pi_{l},\boldsymbol{\mu}_{l},\boldsymbol{\Sigma}_{l},P(c|l)\}_{l=1\dots L,c=1\dots K}$ is the set of model parameters with $\sum_{l=1}^{L}\pi_{l}=1$, $\pi_{l}$ represents the mixing coefficient of the $l$-th gaussian component, $p(\mathbf{f}_{i}|\theta_{l})$ is the probability density function of the $l$-th gaussian component with parameters $\theta_{l}=\{\boldsymbol{\mu}_{l},\boldsymbol{\Sigma}_{l}\}$, and $P(c_{i}|l)$ represents the conditional probability that a feature belongs to class $c_{i}$ given gaussian component $l$.
@@
$\boldsymbol{\mu}_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\mathbf{f}_{i}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\mathbf{f}_{i}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (8)
$\boldsymbol{\Sigma}_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}M_{il}^{(t)}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}M_{il}^{(t)}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (9)
where $M_{il}^{(t)}=(\mathbf{f}_{i}-\boldsymbol{\mu}_{l}^{(t)})(\mathbf{f}_{i}-\boldsymbol{\mu}_{l}^{(t)})^{\top}$.
@@
The model learns its optimal parameters, $\Theta$, by iterating between the Expectation and Maximization until the algorithm converges. Figure 3 offers a visual illustration of the final learned state, presenting a classifier made up of three gaussian components.